### v1

@@ 3   Model Design
For example, a school might aim to assess whether an ADS improves the “academic success” of post-secondary students. A hospital system might evaluate the “quality of care” offered to patients. Yet these concepts are latent constructs that are not directly observable in data (Jacobs and Wallach, 2021). Instead, organizations use readily available proxy outcomes, denoted $\tilde{Y}$ in Figure 1. These are stand-in variables such as post-secondary graduation (for academic success) or hospital readmission (for quality of care). A proxy-construct mismatch arises when such proxy outcomes only imperfectly capture the underlying target construct (Obermeyer et al., 2019; Mullainathan and Obermeyer, 2021; Neil and Zanger-Tishler, 2024), affecting the evaluation science we discuss in Section 4.
@@
Choosing between different problem formulations also depends on domain knowledge. The potential benefits of data-driven decisions can depend on the inherent ease or difficulty of social prediction, the presence vs. absence of heterogeneous causal effects, or both in the particular problem domain. Empirical evidence, either from other contexts or pilot studies, as well as domain knowledge can guide these decisions.
Regarding the difficulty of predicting social outcomes, an important question is how much can complex prediction improve upon simple prediction?
@@
Multiple recent examples echo that simple predictions are competitive with complex machine learning for noisy social outcomes. Simple predictions should be standard benchmarks since they can be implemented for much lower total-cost-of-ownership. Recent studies find that much simpler predictors incorporating only neighborhood/district-level information (Perdomo et al., 2020) or single-variable predictions (Stoddard et al., 2024) are competitive with more complex prediction.141414(Perdomo et al., 2023) find that school and district-level information can target dropout interventions just as effectively as individual risk scores, reflecting the role of structural forces against the variability of individual variation. Stoddard et al. (2024) develop a predictive risk model to predict police misconduct, and that an exceedingly simple model – number of past predictions – is competitive (25%-30% relatively lower recall of top 5% officers by estimated risk) with modern machine learning. While ML culture may favor complex models, this can lead to “worsenalization” (Suriyakumar et al., 2023), when additional covariates worsen predictive performance.
The question of “null efficacy” is also important for causal interventions. Establishing that treatments effects are heterogeneous before considering targeting is necessary. But, a null average treatment effect could have opposite implications as to whether the intervention is overall ineffective, or whether there is underlying heterogeneity that was masked by averaging over subpopulations. It’s folklore that in social domains, treatment effects are often noisy zeroes. It’s harder to test for effect heterogeneity itself is harder than for a null average effect (Crump et al., 2008; Levy et al., 2021a; Sanchez-Becerra, 2023). A mix of empirical and domain-driven justification is necessary.
Characterizing the ceiling of gains from targeting based on covariates is crucial. As always, there are other options, such as expanding the extent of resources given, rather than relying on ADS to mete out scarce resources. Perdomo (2023) theoretically study the value of prediction relative to adding more resources and find that in highly resource-constrained regimes, the impacts of adding more resources typically outweighs the marginal gains of targeting the same scarce resources better.
@@
Even where pure prediction is enough, the ADS context in the social world requires greater care in the design of prediction tasks. The interventional framework can be applied to improve prediction models, in a line of research sometimes called “causal inference for machine learning”.
@@
First of all, the embeddedness of our analyses in the social world, and consequent social dynamics, can affect the validity of model formulations. We discuss in more detail in Appendix B. [^15] To summarize, social interactions complicate predictive validity, for example via strategic, equilibrium or adversarial responses. Though these can be studied theoretically, it’s important to consider fidelity of theoretical assumptions to real-world behavior, in order to avoid inadvertently reinforcing social inequalities by positing rational behavior as a standard. Social dynamics can also affect the validity of the causal inference framework most obviously via interference (outcomes affected by other individuals’ treatment assignments), but in more subtle ways via the inherent conservativity of requiring well-defined (Stevenson, 2023) or modular (Hu, 2023) interventions. Next, even if pure prediction is sufficient, interventional analysis can apply in different ways, which we discuss below.
Causal determinants of outcomes can guide future interventions
Investigating the causal determinants of an outcome can surface new intervention to improve outcomes. Causal determinants are factors that, when changed by some intervention, would change the distribution of the outcome. These may be “direct causes” or “indirect causes” of the outcome, relative to some chosen set of variables. When causal determinants are changeable, they can expand the slate of decision options to improve outcomes. For example, Lee et al. (2023) studied the causal determinants of length of hospital stay after cardiac surgery. Clinicians were interested in which factors — such as medical staffing in the operating room — could be changed to reduce length of stay. Importantly, causal determinants are not the same as the “best predictors” in general. Conversely, the best predictors may not be the best targets for intervention to improve outcomes. Highly predictive variables may be associated with outcomes for multiple reasons, including confounding.
Causal structure of data bias can guide variable selection
Interventional analysis can express complex patterns that shed light on whether or not certain variables should be included in a model. One example is the question of whether protected attribute variables (such as race ) should be included in a predictive model. Some focus on the value that the addition of these variables may have for predictive accuracy (Manski, 2023) depending on the use case (Basu, 2023; Coots et al., 2023) or for improving disparities in realized decisions. Typically, domain knowledge suggests that there is no (or limited) causal direct effect of a protected attribute alone on the downstream outcomes of interest, and the predictive information about race must be due to its correlation with many other social determinants. Causal graphical analysis can precisely describe complex correlation structures and indicate when it makes sense to include information about protected attributes. For example, it may be better to include information on ethnicity when there is label bias (predicting a proxy variable for the true variable of interest) and race has different conditional correlations with the true and proxy variables (Zanger-Tishler et al., 2024).
@@
Interventional analysis, in terms of causal graphical models on covariates, can sharpen debates or highlight finer-grained desiderata. One recent important use was in algorithmic fairness. Researchers have employed causal modeling to articulate how structural or systemic sources of unfairness and injustice surface in the data. More broadly, some causal dependencies among variables used in statistical analyses, including race and gender, might reflect historical and ongoing unfairness at the institutional or systemic level (in addition to inter-personal). Path-specific effects on causal graphs represent hypothesized data-generating processes. Path-specific effects important for studying structural injustices might include the direct effects of race or gender on decisions or outcomes (Kilbertus et al., 2017a; Kusner et al., 2017; Nabi and Shpitser, 2018; Nabi et al., 2022). Researchers can analyze the absence or presence of certain causal effects to argue for context-specific “impermissible” causal dependencies.
Conceptualizing race and gender as causal variables has been controversial: for example, it’s not obvious how to interpret counterfactuals (or causal effects) involving racial categories. Scholars have pointed out that these conceptual difficulties in defining causal effects of race imply that different choices of statistical analysis also correspond to different normative positions on race and the modeling of the social world (Hu and Kohler-Hausmann, 2020; Hu, 2023). Some work in causality-informed algorithmic fairness has suggested that causal definitions of discrimination can adjudicate what differences constitute discrimination on more formal grounds (Kilbertus et al., 2017b). Other work has emphasized that among the potential benefits of using the causal formalism is the incentive to make ethical or normative (as well as empirical) commitments explicit and seek input from affected stakeholders, ethicists, and policymakers (Nabi et al., 2019).
Causal diagrams can clearly delineate different scenarios that parse what is within and outside someone’s control. Imposing desiderata based on path-specific effects might better align predictive models with human moral intuitions and expectations for how decisions should be made. Indeed, (Kiviat, 2019) documents how policymakers’ disagreeements on the moral appropriateness of credit-based insurance scores map onto disagreements about the nature of unobserved confounders that explain observed correlations. [^16] Credit-based insurance scores use financial FICO credit scores to predict how likely someone is to submit an auto-insurance claim. Credit scores were correlated with increased likelihood of submitting an auto-insurance claim. Industry was cagey about the relationship with accidents. While industry argued this correlation was ultimately explained by an unobserved confounder of personal irresponsibility, policymakers expressed concerns that instead it was explained away by income status. It could be that lower-income individuals had lower credit score, and also higher rates of making claims vs. higher-income individuals at the same rate of accidents. Using causal structure to clearly delineate terms of current disagreement can guide future deliberation.
@@ 4   Evaluation Science
Following the model design of predictive ADS, an equally important consideration is how to evaluate whether these systems are working (or will work) as intended. In the context of ADS, evaluation science is the scientific process of quantifying the gap between articulated expectations for a system and the measured reality of that system’s behavior or output  (Raji, 2022; Mathison, 2004). Effective evaluation of predictive ADS systems must involve studying whether their use and prediction aligns with their stated intent and underlying policy objectives, while balancing tradeoffs in resource allocation, e.g., between pre- and post-deployment assessments.
There are multiple evaluation approaches that could provide meaningful evidence to assess common claims about ADS. A widely used approach for performance measurement is benchmarking (Section 4), which evaluates the predictive statistics of the model underlying the ADS, such as accuracy, ROC scores, or F1 metrics on curated datasets, against a chosen proxy outcome. However, while useful, benchmarking provides only partial evidence for the performance of ADS, as high predictive accuracy alone rarely meets the real-world objectives of ADS deployments (Liu et al., 2023a; Perdomo et al., 2023). There may also be concerns with the validity of the evaluation (Section 4) due to measurement challenges such as label bias and proxy-construct mismatch.
This motivates an interventional evaluation paradigm that moves beyond benchmarking to consider how predictions affect decisions and outcomes in real world deployments. In Section 4, we outline such a paradigm for assessing the causal impacts of predictive ADS on outcomes of interest, drawing upon recent work such as Imai et al. (2023), while also highlighting open research challenges such as off-policy evaluation, the validity of causal inference, and the complexities of experimental design involving interference and human decision makers. Finally, in Section 4, we survey the broader social and organizational challenges in evaluating ADS, including the challenge of clearly articulating the goals and standards for an ADS system itself.
@@
Benchmarking fails for ADS
A critical part of the evaluation process is operationalizing policy goals into one or more target outcomes ($Y$) to use for evaluation. To set the stage for a new framework of interventional evaluation, we overview the standard paradigm of benchmarking.
Limitations of Benchmarking
Standard benchmarking seeks to characterize the prediction error of the ADS (e.g. a risk score denoted $R$), against a target $Y$, using a static dataset. Predictive models might be compared on an alphabet soup of different error metrics. Evaluations of predictive models based on these large shared benchmark datasets typically follow a classical train-test split. Donoho (2024) attributes the success of data science to the “frictionless reproducibility” of this exact benchmark-driven culture: shareable and open datasets, development of open-source and reproducible methodology, and competitive testing on shared benchmarks.
@@
So far, the evidence for and against the use of such risk assessments illustrates the broader dialogue on how to assess the impact of algorithmic interventions in high-stakes settings. This scenario thus provides an ideal case study in examining the different approaches to evaluation in this context, prompting a discussion on limitations to these approaches, as well as opportunities for methodological improvement.
@@
Validity concerns are pervasive in both benchmarking and interventional evaluations. The concept of validity describes the extent to which an evaluation performed via one or more proxy outcomes captures the intended properties of a system (Coston et al., 2023).
Assuming that ideal outcomes can be defined as prediction targets, it is necessary to evaluate the measurement error and validity of benchmarking. A common reason why validity can fail in practice is that the evaluation is conducted under conditions which do not represent real-world deployment. This threatens external validity. We go into more detail about two pervasive issues: label bias, when available outcomes are proxies of ideal outcomes for evaluation, and selection bias, resulting in available data that systematically differs from the full population.
@@
Label bias can arise from practical compromises when choosing outcomes in problem formulation.
@@
Proxy outcomes used to evaluate an ADS may imperfectly reflect broader policy goals. This measurement gap raises key evaluation challenges: How do we select the most relevant proxy outcomes to track? How do we characterize potential gaps between proxy outcomes and ideal targets? Should we attempt to collect additional data? The following discussion explores these challenges and suggests approaches to ensure evaluations measure progress toward intended policy goals.
@@
In the case of arrest and offending, this problem has long been known in criminology (Biderman and Reiss Jr, 1967), and scholars have suggested various ways to think about dealing with this data mismatch (Zilka et al., 2023; Neil and Zanger-Tishler, 2024; Nagaraj et al., 2025; Schoeffer et al., 2025).
Several strategies provide evidence for the validity of evaluations performed using proxy outcomes. First, showing that the conclusions of an RAI evaluation hold across multiple proxy outcomes (e.g., re-arrest, failure to appear) provides evidence for its validity (Kleinberg et al., 2018; De-Arteaga et al., 2018). Second, demonstrating that an RAI evaluation holds under different plausible assumptions on the relationship between target and proxy outcomes (i.e., sensitivity analysis (Fogliato et al., 2020)) provides evidence for validity. Additional research to develop tools to help practitioners to assess validity would be helpful.
The validity of proxy outcomes can break down in two key ways: sometimes the target construct is fundamentally unobservable (like true criminal behavior (Zilka et al., 2023; Neil and Zanger-Tishler, 2024; Biderman and Reiss Jr, 1967)), while in other cases the target construct could be measured but practical barriers prevent its collection (like detailed health needs assessments). Understanding the type of measurement challenge is essential for determining how to ensure valid RAI evaluation (Mullainathan and Obermeyer, 2021).
New data collection methods may be needed to accurately assess RAIs’ effects on the actual outcomes of interest. Evaluations of RAIs should take into account evaluating progress on longer-term policy goals that motivate their deployment. An RAI might achieve short-term goals of reducing detention. Studying effects on crime rates can bolster the evidence base. Traditional data sources often fail to capture these nuanced effects, making innovative data collection approaches critical. But data collection comes at a cost, including important social risks and ethical concerns, especially in sensitive domains like criminal justice or child welfare. Additional surveillance of decision subjects could raise serious privacy and consent issues.
@@
We illustrate the challenges of selective labels in the example of bail decisions (the key empirical example in (Lakkaraju et al., 2017)). Bail decisions have a somewhat complex relationship with the pretrial detention task discussed in Example 1. A judge makes the decision to set bail or not, and if so; what amount. In turn, the defendant posts (pays) bail (if it is a secured bond), or not. If an individual cannot post bail, in the sense that they cannot arrange funds to pay the secured bond, they may be detained pretrial. If an individual is detained pretrial $(Z=1)$, they definitionally cannot fail to appear for their trial and their ability to commit another crime is severely curtailed. If an individual makes bail, they are released into the community (albeit with a financial incentive to appear for trial), and all outcomes are possible: they may make their trial or commit another crime.
@@
Therefore, prior bail decisions censor the outcomes that risk-assessment instruments are designed to predict. Now if an individual is detained pretrial ($Z=1$), their downstream outcomes are censored. From the perspective of the individual who was detained pretrial, if they had in fact not been detained, there are potential outcomes of whether they would have failed or succeeded to appear for trial (FTA), and whether or not they would have committed another crime (NCA/NVCA); the exact outcomes RAIs are designed to predict. The latter is a counterfactual perspective. An alternative approach to formulate evaluations is a consequentialist perspective that assigns utility or cost to final outcomes (failure to appear or committing another crime) and decisions (societal cost of bail/detention) marginally. Unfortunately, the counterfactual perspective depends on joint potential outcomes (what would have happened if an individual was not in fact detained), and is unidentifiable without assumptions like monotonicity. However, many policy-relevant arguments against or for cash bail exactly do reflect this individual counterfactual perspective, whether or not cash bail “punishes the poor” or otherwise improves outcomes via “specific deterrence”. [^17] More specifically, these arguments posit that cash bail “punishes the poor” by causing detention for poor individuals unable to afford bail bond who otherwise would not have offended. Or, they argue on the contrary that money bail provides “specific deterrence” in that it provides financial incentives for an individual to avoid failure to appear relative to if they had in fact been released on recognizance, without any conditions.
@@
The problem of selective labels is pervasive beyond criminal justice alone. Prior decisions $(Z=1)$ censor many outcomes of interest, especially those that screen in or out individuals. For example, banks do not observe loan default or payback outcomes for individuals they deny loans for. [^18] In lending, they study methods for addressing this under the terminology of “reject inference” (Banasik and Crook, 2007; Ehrhardt et al., 2021). Therefore benchmarking evaluations on available data are assessing, for example, error of a predictor $\mathbb{E}[(Y-\mu(X))^{2}\mid Z=1]$ on the censored-in population. This selection bias poses a key challenge for external validity: researchers found that evaluations of credit-lending algorithms on a sample of approved applicants systematically misrepresents the accuracy and fairness properties on the full population of loan applicants due to distribution shifts (Coston et al., 2021; Chien et al., 2023).
@@
This section overviews approaches that evaluate the causal effect of predictive ADS ($R$) on outcomes of interest ($Y$). Different evaluation approaches are available and appropriate depending on experimentation capacity and information. Differences in experimentation capacity include whether only historical data can be used for quasi-experimental and observational evaluation, or whether randomizing ADS access is possible. Differences in information include whether only organization-level information on deployment of ADS and aggregate outcomes is available, or individual-level information. We develop a toolbox for interventional analysis based on breaking down the complex interactions among predictive risk scores, actual decisions (possibly made by human decision-makers), and outcomes, into simpler analyses between pairs therein.
Navigating the hierarchy of evidence: interventional evaluation can meet organizations where they are
Randomized controlled trials are often considered the gold standard of evidence of causal impacts. Because the status quo decision-making process influences outcomes of interest (e.g., health status, student success), it is challenging to compare alternative decision-making processes without experimental assignment of policy alternatives. Yet, experimentation is both costly, resource intensive, and ethically fraught depending on the proposed intervention.
Sometimes ADS program evaluation takes place only after a system has been deployed. But the effects of ADS can also be estimated prior to deployment using retrospective observational data for off-policy evaluation. In practice, rigorous policy evaluation may require a combination of both pre-deployment evaluation via retrospective observational data and post-deployment evaluation following experimental assignment of the policy alternatives.
Figure 6: Overview of different interventional evaluation paradigms. Section 4 evaluates impacts of changes in decisions ($D$) on outcomes ($Y$), i.e. ITR performance evaluation, highlighted in blue. Some quasi-experimental approaches consider the effect of introducing ADS on outcomes $Y$, highlighted in green. Other analyses (observational or randomized) use micro-level data to investigate interactions between risk scores $R$ and decisions $D$, and $Y$; highlighted in red.
There are different ways of operationalizing the effects of ADS on outcomes. Often ADS impacts downstream outcomes by informing specific decisions or interventions (denoted $D$). We begin by overviewing evaluation of differences in $D$ on $Y$, aka evaluating ITRs that assume direct intervention on decisions received.
Additionally accounting for the complexity of how ADS ($R$) affects $D$ and therefore $Y$ requires some additional modeling; there are more degrees of freedom. We discuss different options, based on the types of data available. Without individual-level data on decisions or covariates, evaluation requires conceptualizing the provision of $R$ as a policy intervention. With micro-level data on $(R,D,Y(D))$, the simplest case is if the decision $D$ is a function of the risk score; if not, then this requires additional modeling of how risk scores inform decisions. We discuss quasi-experimental and observational causal inference approaches in this setting.

### v2

@@ 3  Model Design
Regarding the difficulty of predicting social outcomes, an important question is how much can complex prediction improve upon simple prediction? A mix of domain and empirical knowledge is required to answer this question, and social domains are often categorically different than other engineering “signal recovery” domains.
@@
Multiple recent examples echo that simple predictions are competitive with complex machine learning for noisy social outcomes. Simple predictions should be standard benchmarks since they can be implemented for much lower total-cost-of-ownership. Recent studies find that much simpler predictors incorporating only neighborhood/district-level information (Perdomo et al., 2023) or single-variable predictions (Stoddard et al., 2024) are competitive with more complex prediction.171717(Perdomo et al., 2023) find that school and district-level information can target dropout interventions just as effectively as individual risk scores, reflecting the role of structural forces against the variability of individual variation. Stoddard et al. (2024) develop a predictive risk model to predict police misconduct, and that an exceedingly simple model – number of past predictions – is competitive (25%-30% relatively lower recall of top 5% officers by estimated risk) with modern machine learning. While ML culture may favor complex models, this can lead to “worsenalization” (Suriyakumar et al., 2023), when additional covariates worsen predictive performance.
The question of “null efficacy” is also important for causal interventions. Establishing that treatments effects are heterogeneous before considering targeting is necessary. But, a null average treatment effect could have opposite implications as to whether the intervention is overall ineffective, or whether there is underlying heterogeneity that was masked by averaging over subpopulations. It’s folklore that in social domains, treatment effects are often noisy zeroes. It’s harder to test for effect heterogeneity itself is harder than for a null average effect (Crump et al., 2008; Levy et al., 2021a; Sanchez-Becerra, 2023).
Contextualizing the Relative Value of Prediction in Resource Allocation Problems
Given that prediction is just one component of a broader system, there are many other ways of driving improvements in downstream social welfare apart from optimizing prediction quality. For example, directly expanding resource capacity and assigning interventions to more people, or just investing in increasing the effect sizes of the interventions themselves.
Perdomo (2024) and Fischer-Abaigar et al. (2025) study how do the welfare benefits of investing in prediction compare to those of other policy levers such as adding more resources. Through theoretical analysis and an in-depth case studied on unemployment prediction in Germany, they find that in highly resource-constrained regimes the impacts of adding more resources typically outweighs the marginal gains of targeting the same scarce resources better.
@@
Even where pure prediction is enough, the ADS context in the social world requires greater care in the design of prediction tasks. The interventional framework can be applied to improve prediction models, in a line of research sometimes called “causal inference for machine learning”. First of all, the embeddedness of our analyses in the social world, and consequent social dynamics, can affect the validity of model formulations. We discuss in more detail in Appendix B. [^18] To summarize, social interactions complicate predictive validity, for example via strategic, equilibrium or adversarial responses. Though these can be studied theoretically, it’s important to consider fidelity of theoretical assumptions to real-world behavior, in order to avoid inadvertently reinforcing social inequalities by positing rational behavior as a standard. Social dynamics can also affect the validity of the causal inference framework most obviously via interference (outcomes affected by other individuals’ treatment assignments), but in more subtle ways via the inherent conservativity of requiring well-defined (Stevenson, 2023) or modular (Hu, 2023) interventions. Next, even if pure prediction is sufficient, interventional analysis can apply in different ways, which we discuss below.
Causal structure of data bias can inform variable selection Interventional analysis can express complex patterns that inform deliberations of including covariate information in a predictive model (or decision rule) or not. One example is the question of whether protected attribute variables (such as race) should be included in a predictive model. Some focus on the value that the addition of these variables may have for predictive accuracy (Manski, 2023) depending on the use case (Basu, 2023; Coots et al., 2023) or for improving disparities in realized decisions. Insofar as race is understood to be a social category marking social difference (rather than, say, biological difference), the reasons that underlie race’s predictive power in various social systems largely trace to those systematic social differences that define a racialized social structure. A world in which important life outcomes are racially stratified is a world in which race has powerful predictive power. The reasons for racial stratification today are of course the legacy of decades and centuries of race-making processes which differentially distribute(d) power, resources, rights, esteem, among other benefits and burdens along lines of perceived geographic ancestry or phenotype that we now call ”racial”. Causal graphical analysis, alongside other historical and social-scientific analyses, can be used to help draw out these complex correlation structures.
@@
Interventional analysis, in terms of causal graphical models on covariates, can sharpen debates or highlight finer-grained desiderata. One recent important use was in algorithmic fairness. Researchers have employed causal modeling to articulate how structural or systemic sources of unfairness and injustice surface in the data. More broadly, some causal dependencies among variables used in statistical analyses, including race and gender, might reflect historical and ongoing unfairness at the institutional or systemic level (in addition to inter-personal). By constructing simplified causal models of complex social systems, one can trace schematically “path-specific” causal effects representing various causal processes that generate the data we observe today. Path-specific effects can be fruitfully used to identify particular processes that generate structural injustices along race or gender lines (Kilbertus et al., 2017a; Kusner et al., 2017; Nabi and Shpitser, 2018; Nabi et al., 2022). Concrete claims as to how certain race or sex causal effects are generated can help facilitate broader moral and political arguments for why certain causal dependencies should or should not be leveraged in prediction.
That said, conceptualizing race and gender as causal variables has been controversial: for example, it’s not obvious how to interpret counterfactuals (or causal effects) involving racial categories. Scholars have argued that these conceptual difficulties in defining causal effects of race imply different normative positions on what race is and how the social world works(Hu and Kohler-Hausmann, 2020; Hu, 2023). Other work in causality-informed algorithmic fairness has suggested that causal definitions of discrimination can help to adjudicate what differences constitute discrimination on more formal grounds and hopefully clearer (Kilbertus et al., 2017b). Other work has emphasized that among the potential benefits of using the causal formalism is the incentive to make ethical or normative (as well as empirical) commitments explicit and seek input from affected stakeholders, ethicists, and policymakers (Nabi et al., 2019).
@@ 4  Evaluation Science
Following the model design of predictive ADS, an equally important consideration is how to evaluate whether these systems are working (or will work) as intended. In the context of ADS, evaluation science is the scientific process of quantifying the gap between articulated expectations for a system and the measured reality of that system’s behavior or output  (Raji, 2022; Mathison, 2004).
Multiple evaluation approaches provide meaningful evidence to assess claims about ADS. A widely used approach for performance measurement is benchmarking (Section 4), which evaluates the predictive statistics of the model underlying the ADS, such as accuracy, ROC scores, or F1 metrics on curated datasets, against a chosen proxy outcome. However, while useful, benchmarking provides only partial evidence for the performance of ADS, as high predictive accuracy alone rarely meets the real-world objectives of ADS deployments (Liu et al., 2023a; Perdomo et al., 2023). There may also be concerns with the validity of the evaluation (Section 4) due to measurement challenges such as label bias and proxy-construct mismatch.
This motivates an interventional evaluation paradigm that moves beyond benchmarking to consider how predictions affect decisions and outcomes in real world deployments. In Section 4, we highlight different options for evaluating along the causal chain of $X\to R\to D\to Y$ as well as novel causal challenges that arise in doing so. Finally, in Section 4, we survey the broader social and organizational challenges in evaluating ADS, including the challenge of clearly articulating the goals and standards for an ADS system itself.
@@
Benchmarking Failures for ADS
A critical part of the evaluation process is operationalizing policy goals into one or more target outcomes ($Y$) to use for evaluation. To set the stage for a new framework of interventional evaluation, we overview the standard paradigm of benchmarking. Standard benchmarking seeks to characterize the prediction error of the ADS (e.g. a risk score denoted $R$), against a target $Y$, using a static dataset. Predictive models might be compared on an alphabet soup of different error metrics. Donoho (2024) attributes the success of data science to the “frictionless reproducibility” of this exact benchmark-driven culture: shareable and open datasets, development of open-source and reproducible methodology, and competitive testing on shared benchmarks.
@@
So far, the evidence for and against the use of such risk assessments illustrates the broader dialogue on how to assess the impact of algorithmic interventions in high-stakes settings. We continue with case studies that highlight different approaches to evaluation, even for the same ADS and policy context.
@@
Validity concerns are pervasive in both benchmarking and interventional evaluations. The concept of validity describes the extent to which an evaluation performed via one or more proxy outcomes captures the intended properties of a system (Coston et al., 2023). A common reason why validity can fail in practice is that the evaluation is conducted under conditions which do not represent real-world deployment. This threatens external validity. We go into more detail about two pervasive issues: label bias, when available outcomes are proxies of ideal outcomes for evaluation, and selection bias, resulting in available data that systematically differs from the full population.
@@
Label bias can arise from practical compromises when choosing outcomes in problem formulation. For example, a school might aim to assess whether an ADS improves the “academic success” of post-secondary students. A hospital system might evaluate the “quality of care” offered to patients. Yet these concepts are latent constructs that are not directly observable in data (Jacobs and Wallach, 2021). Instead, organizations use readily available proxy outcomes, denoted $\tilde{Y}$ in Figure 1. These are stand-in variables such as post-secondary graduation (for academic success) or hospital readmission (for quality of care). A proxy-construct mismatch arises when such proxy outcomes only imperfectly capture the underlying target construct (Obermeyer et al., 2019; Mullainathan and Obermeyer, 2021; Neil and Zanger-Tishler, 2024).
@@
Proxy outcomes used to evaluate an ADS may imperfectly reflect broader policy goals. This measurement gap raises key evaluation challenges: How do we select the most relevant proxy outcomes to track? How do we characterize potential gaps between proxy outcomes and ideal targets? Should we attempt to collect additional data?
@@
In the case of arrest and offending, this problem has long been known in criminology (Biderman and Reiss Jr, 1967), and scholars have suggested various ways to think about dealing with this data mismatch (Zilka et al., 2023; Neil and Zanger-Tishler, 2024; Nagaraj et al., 2025; Schoeffer et al., 2025). The validity of proxy outcomes can break down in two key ways: sometimes the target construct is fundamentally unobservable (like true criminal behavior (Zilka et al., 2023; Neil and Zanger-Tishler, 2024; Biderman and Reiss Jr, 1967)), while in other cases the target construct could be measured but practical barriers prevent its collection (like detailed health needs assessments).
Several strategies can bolster validity of conducting evaluations with proxy outcomes, such as assessing against multiple proxy outcomes (e.g., re-arrest, failure to appear) (Kleinberg et al., 2018; De-Arteaga et al., 2018). Demonstrating that an RAI evaluation holds under different plausible assumptions on the relationship between target and proxy outcomes (i.e., sensitivity analysis (Fogliato et al., 2020)) provides evidence for validity.
@@
This issue arises in bail decisions, a slight variant of the PSA setup in Example 1. Setting bail is an additional decision in between detention without bail or release without conditions. Individuals who are offered bail may be unable to pay it, and then detained pretrial [^19] The defendant posts (pays) bail (if it is a secured bond), or not. If an individual cannot post bail, in the sense that they cannot arrange funds to pay the secured bond, they may be detained pretrial. If an individual makes bail, they are released into the community (albeit with a financial incentive to appear for trial), and all outcomes are possible: they may make their trial or commit another crime. . If an individual is detained pretrial $(Z=1)$, they definitionally cannot fail to appear for their trial and their ability to commit another crime is severely curtailed; their downstream outcomes are censored. Therefore, changes in bail policy introduce distribution shifts — the New York Criminal Justice Agency’s validation of the release assessment (Kitzmiller and Gewirtz, 2025) finds that average risk of released individuals increased after bail reform resulted in releasing significantly more individuals.
@@
From the perspective of the individual who was detained pretrial, if they were released instead, they may or may not have re-appeared for trial (1-FTA), or committed another crime (NCA). This is a counterfactual that depends on joint potential outcomes — what would have happened if an individual was released instead. Contrast this with a societal consequentialist perspective that assigns utility or cost to final outcomes (failure to appear or committing another crime) and decisions (societal cost of bail/detention) marginally — viewing detention as equivalent to reducing failure to appear. The individual counterfactual perspective resonates with moral debates whether or not cash bail “punishes the poor” or otherwise improves outcomes via “specific deterrence”. [^20] More specifically, these arguments posit that cash bail “punishes the poor” by causing detention for poor individuals unable to afford bail bond who otherwise would not have offended. Or, they argue on the contrary that money bail provides “specific deterrence” in that it provides financial incentives for an individual to avoid failure to appear relative to if they had in fact been released on recognizance, without any conditions. But, it is not identified from observed data and requires additional assumptions or structure. Nonetheless, the contrast highlights that selective labels censor valuable information and introduce distribution shift challenges.
@@
The problem of selective labels is pervasive beyond criminal justice alone. For example, banks do not observe loan default or payback outcomes for individuals they deny loans for. [^21] In lending, they study methods for addressing this under the terminology of “reject inference” (Banasik and Crook, 2007; Ehrhardt et al., 2021). Therefore benchmarking evaluations on available data are assessing, for example, error of a predictor $\mathbb{E}[(Y-\mu(X))^{2}\mid Z=1]$ on the censored-in population [^22] This selection bias poses a key challenge for external validity: researchers found that evaluations of credit-lending algorithms on a sample of approved applicants systematically misrepresents the accuracy and fairness properties on the full population of loan applicants due to distribution shifts (Coston et al., 2021; Chien et al., 2023). . In some settings, institutions can take advantage of randomized samples to avoid selective labels. The IRS, for instance, maintained a “National Research Program,” which is a stratified random sample of taxpayers who are audited, which can help overcome selective labels (Elzayn et al., 2024). But the presence of such representative samples with labels of interest can be rare in resource constrained settings.
@@
We now turn from benchmarking to interventional evaluation: studying how predictive systems causally affect outcomes. Formally, an ADS produces a risk score $R$ that influences decisions $D$, which in turn affect outcomes $Y$. Different empirical settings reveal different parts of this chain. The appropriate evaluation design depends on both (i) experimentation capacity—whether randomized assignment of decisions or ADS access is possible—and (ii) information availability—whether we observe individual-level $(R,D,Y)$ or only aggregate outcomes.
Hierarchy of evidence.
Randomized controlled trials remain the most credible way to establish causal effects, yet they are often infeasible in high-stakes institutional settings. Rigorous evaluation therefore meets organizations “where they are,” combining designs based on retrospective observational data, natural or quasi-experiments, and prospective randomized rollouts. In practice, robust evidence may require both pre-deployment (off-policy or sandbox) analyses and post-deployment field trials.
Figure 6: Our interventional evaluation paradigm vs. the status quo of “blind men studying the elephant”, highlighted. Blue: $D\to Y$ (ITR evaluation). Green: policy-level impacts of introducing an ADS (event studies). Red: $R\to D\to Y$ with micro-level data (human–algorithm interaction and encouragement designs).
Our integrated framework: $X\to R\to D\to Y$.
We emphasize the causal chains between predictions, interventions and outcomes. To organize the growing literature, we decompose evaluation efforts by the links they estimate. We begin by overviewing methods that assess the causal impact of decisions or interventions on outcomes $D\!\to\!Y$, such as evaluating an individualized treatment rule or targeted policy. Other approaches for the effects of predictions appeal to the presence of some policy change that introduces ADS: based on macro-data alone, some approaches estimate aggregate effects of introducing or withdrawing ADS (e.g., via event studies or difference-in-differences). Elsewhere, other studies focus on $R\!\to\!D$: studying how human decision-makers rely on or respond to risk scores. One case that simplifies estimating the effect of $R\to D\to Y$ is if the decision $D$ is a function of the risk score; else, there are fundamental causal challenges. We overview different approaches. Altogether, viewing these disparate perspectives in our integrated framework provides a unified toolbox for understanding the impacts of ADS on social welfare.